OpenStudy (tylerd):

dy/dx (cot(5+7x)(7+6x))

OpenStudy (tylerd):

i know the answer but im having trouble with a certain part of it

OpenStudy (dumbcow):

what is inside the cot function ?

OpenStudy (tylerd):

cot(5+7x)

OpenStudy (tylerd):

csc2(5+7x) (−49−42x) + 6cot(5+7x) not sure where the -49-42x comes from

OpenStudy (tylerd):

dunno why that post is all messed up for me?

OpenStudy (dumbcow):

ok just checking you would use both product rule and chain rule product rule part: \[6 \cot(5+7x) + [\cot(5+7x)]' (7+6x)\] chain rule part \[[\cot(5+7x)]' = -\csc^2 (5+7x) * (5+7x)' = -7 \csc^2(5+7x)\]

OpenStudy (dumbcow):

i think the -49-42x is from distributing the -7 ---> -7(7+6x) = -49 -42x

OpenStudy (dumbcow):

@TylerD , try refreshing your browser, it may make the ??? go away in your post

OpenStudy (tylerd):

ok gonna be a few mins before i make sense of that one sec.

OpenStudy (tylerd):

thanks